OpenStudy (anonymous):

Medal rewarded!! Just one quick question please help Given the trinomial, what is the value of the coefficient 'B' in the factored form? 2x2 + 14xy - 36y2 = 2(x + By)(x - 2y) B= ____________

OpenStudy (anonymous):

2*-36=-72 18-4=14 18*-4=-72 write 14 as (18-4) and make factors of left hand side.

OpenStudy (anonymous):

im so confused @surjithayer

OpenStudy (anonymous):

\[2 x^2+14 xy-36 y^2=2\left( x^2+7 x y-18 y^2 \right)=2\left\{ x^2+\left( 9-2 \right)xy-18y^2 \right\}\] \[=2\left\{ x^2+9xy-2xy-18 y^2 \right\}=2\left\{ x \left( x+9 y \right)-2 y \left( x+9 y \right) \right\}\] \[=2\left( x+9 y \right)\left( x-2 y \right)\] B=?

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

correct.